OpenStudy (anonymous):

please comrades can someone clearly show me steps in arriving at the derivative of the logarithmic function y = ln x which is 1/x??

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

(ln(x+h)- ln (x))/h ln (1 + h/x) / h = ln(1 + h/x) / hx/x = 1/x (ln(1+h/x) / h/x)

OpenStudy (anonymous):

the log part tends to 1 as h goes to zero..standard definition limit